OpenStudy (tl155):

103 • [12 ÷ (10 – 4 • 2)] + 3

OpenStudy (welshfella):

this is an exercise in order of operations. have you heard of PEMDAS?

OpenStudy (tl155):

yes but i have a lot of trouble

OpenStudy (welshfella):

well work out the innermost parentheses first 10 - 4*2 = do the multiplication first

OpenStudy (welshfella):

* i meant - not =

OpenStudy (sunrainwolf):

Work the brackets first, PEMDAS, Parenthese, exponents, multiplication, division, then adding then finally subtraction.

OpenStudy (welshfella):

10 - 4*2 multiplication comes before subtract so do the 4 * 2 first then subtract from 10

OpenStudy (welshfella):

what do you get?

OpenStudy (tl155):

is the answer 621

OpenStudy (welshfella):

can you show me your work?

OpenStudy (tl155):

\[4x2=8-10=2x103=206+3=209\]

OpenStudy (tl155):

so 209

OpenStudy (sunrainwolf):

103 * [12/(10-4*2)]+3 103*[12/(10-8)]+3 103*[12/2]+3 103*6+3 618+3 621 When I worked the problem this is what I got. You work work 4*2 first in the brackets, 10-8 is 2, 12 divided by 2 is 6, 103 times 6 is 613 then add 3 to make it 621.

OpenStudy (tl155):

ok